If 9.0 of oxygen gas are consumed with excess propane in the following reaction, how many miles of carbon dioxide will be produced?

C3H8(g)+5O2(g)= 3CO2(g)+4H2O(g)
1
Expert's answer
2020-03-30T10:50:14-0400

m (O2) = 9.0 g

M = 32.0 g/mol

n = 9.0 / 32.0 = 0.28125 mol


n (CO2) = 3/5 * 0.28125 = 0.16875 mol

V = 0.16875 * 22.4 L/mol = 3.78 L = 3,780 mL


if 9.0 is volume, for example 9.0 mL, then 9.0 * 3/5 = 5.4 mL CO2 is produced.